​A POINT MOVES IN A STRAIGHT LINE SO THAT ITS DISPLACEMENT ‘x’( in meter) AT TIME ‘t’(in seconds) IS GIVEN BY -
‘ x= t2 + 1. ITS ACCELERATION IN m/sec2, at time ‘t’ is-
  1. 1/x3
  2. 1/x – 1/x2
  3. -1/x2
  4. -( t2/x3)

Grade:Select Grade

4 Answers

VISHESH DHARAIYA
9 years ago
x2=t2+1
taking its differentiation,
2x*(dx/dt)=2t
x*(dx/dt)=t
again taking differentiation for acceleration,
x*(d2x/dt2) + (dx/dt)2 = 1
x*(d2x/dt2) = 1 – (dx/dt)2
x*(d2x/dt2) = 1 – (t/x)2                  (because x*(dx/dt)=t)
(d2x/dt2) = (1 – (t/x)2)/x
(d2x/dt2) = 1/x – t2/x3
 
put t2 = x2 -1
(d2x/dt2) = 1/x3
 
 
 
Bharat Makkar
9 years ago
thanks!!!!
Pooja
4 years ago
 
x2=t2+1
taking its differentiation,
2x*(dx/dt)=2t
x*(dx/dt)=t
again taking differentiation for acceleration,
x*(d2x/dt2) + (dx/dt)2 = 1
x*(d2x/dt2) = 1 – (dx/dt)2
x*(d2x/dt2) = 1 – (t/x)2                  (because x*(dx/dt)=t)
(d2x/dt2) = (1 – (t/x)2)/x
(d2x/dt2) = 1/x – t2/x3
 
put t2 = x2 -1
(d2x/dt2) = 1/x3
 
Hope this will help you
.....
ankit singh
3 years ago
x2=t2+1
taking its differentiation,
2x*(dx/dt)=2t
x*(dx/dt)=t
again taking differentiation for acceleration,
x*(d2x/dt2) + (dx/dt)2 = 1
x*(d2x/dt2) = 1 – (dx/dt)2
x*(d2x/dt2) = 1 – (t/x)2                  (because x*(dx/dt)=t)
(d2x/dt2) = (1 – (t/x)2)/x
(d2x/dt2) = 1/x – t2/x3
 
put t2 = x2 -1
(d2x/dt2) = 1/x3